OpenStudy (anonymous):

Can you please help me!!! A ship is sailing parallel with a straight coastline, but the captain doesn’t know how far from the shore his is. A lighthouse is visible to the ship, and the angle between the line of the ship and the lighthouse is 30◦. After traveling 3 miles, the captain measures the angle to be 75◦. How far does he determine the boat to be from the coastline [decimal preferred]? (This is not the distance from the boat to the lighthouse.)

OpenStudy (sburchette):

I tried this problem but I'm not sure if it is correct or not, its just my jab at it? Do you happen to have the solutions available? I got a distance of about 2 miles. I can upload some diagrams once I draw them up.

OpenStudy (anonymous):

Unfortunatly, I don't have the solutions available at this time. But, I would like to see your diagrams when they are ready....

OpenStudy (sburchette):

All right, I'm cranking them out as we speak =)

OpenStudy (anonymous):

Awesome! :)

OpenStudy (anonymous):

tan 30 = y/x tan 75 = y/ ( x -3) => √3 x/3 = √ 3 + 2 ( x -3)

OpenStudy (anonymous):

x is the distance at 30 degree! -> x -3 is the distance at 75 degree.

OpenStudy (sburchette):

But if the ship is sailing parallel with the coastline, wouldn't the distance from the shoreline remain constant?

OpenStudy (anonymous):

That's what I thought too....

OpenStudy (anonymous):

From my pic match up with pic you post, that how I come up with the solution.

OpenStudy (anonymous):

My x is y in the pic, ( somehow my vision just can see my way) From x => y which is the parallel distance ( x in the your pic)

OpenStudy (anonymous):

Yes, the distance y ( = x in your pic) is constant!

OpenStudy (sburchette):

Ok, we just set just assigned different variables to the sides.

OpenStudy (anonymous):

Solve from this system equa: tan 30 = y/x tan 75 = y/ ( x -3)

OpenStudy (anonymous):

So that's the answer then?

OpenStudy (anonymous):

All you do is solve it to get y is the parallel distance!

OpenStudy (anonymous):

Solve the above equation, right?

OpenStudy (anonymous):

Yes, tan 30 = √3/3, tan 75 = 3.732

OpenStudy (anonymous):

Ok! :) Thanks guys so much for helping me out!!!